A particular drug is supplied in tablets of 600 mg strength.A 90 lb patient requires a one week supply of this drug at a dose of 20mg/lb every 12 hours. How many tablets are required in total?
Answer #1

A particular drug is supplied in tablets= 600 mg strength

Patient’s weight= 90 lb

The patient requires a one week supply of this drug at a dose of 20mg/lb every 12 hours.

Required dose = 20mg/lb

Means for 1 lb=20 mg

For 90 lb= 20X 90=1800 lb for every 12 hours

Then for one day /everyday/24 hours required dose = 1800X2=3600 mg

Then for one week (7 days) required dose= 3600X 7= 25,200 mg total

600 mg =1 tablet

1 mg=1/600 tablet

25,200 mg= (1/600) X 25200= 42 tablets answer

Hence Total 42 tablets will be required.